Two cards are drawn without replacement from a standard deck of 52 playing cards. What is the probability
of choosing a club and then, without replacement, a black card? Write your answer as a fraction or a decimal
number rounded to three decimal places.
Answer #1

Ans:

There are 26 black cards,out of which 13 are clubs.

Without replacement:

P(a club and a black card)=(13/52)*(25/51)=25/204 or 0.123